# Understanding the Potential of FPGA-Based Spatial Acceleration for Large Language Model Inference

### 23 Dec 2023 v1

@@ Abstract.
To enable more productive implementations of an LLM model on FPGAs, we further provide a library of high-level synthesis (HLS) kernels that are composable and reusable. This library will be made available as open-source. To validate the effectiveness of both our analytical model and HLS library, we have implemented BERT and GPT2 on an AMD Xilinx Alveo U280 FPGA device. Experimental results demonstrate our approach can achieve up to 16.1$\times$ speedup when compared to previous FPGA-based accelerators for the BERT model. For GPT generative inference, we attain a 2.2$\times$ speedup compared to DFX, an FPGA overlay, in the prefill stage, while achieving a 1.9$\times$ speedup and a 5.7$\times$ improvement in energy efficiency compared to the NVIDIA A100 GPU in the decode stage.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: TRETS’24; May 5–8, 2024; Orlando, FLisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
The rapid advancement of Transformer-based large language models (LLMs) (Vaswani et al., 2017; Bommasani et al., 2021) has sparked a revolution across a wide range of natural language processing tasks, such as conversational AI (OpenAI, 2023; Zheng et al., 2023; Cohen et al., 2022) and code generation (Chen et al., 2021; Li et al., 2022; Nijkamp et al., 2023). Recent research has brought to light the phenomenon of “emergence” in LLMs, where advanced capabilities become evident as the models scale up to billions of parameters (Wei et al., 2022a, b). However, supporting this unprecedented scale poses significant challenges, particularly in terms of computational and memory resources. At the same time, the increasing use of LLMs in interactive applications like voice assistants and autonomous systems requires hardware accelerators capable of providing both low latency and high energy efficiency (OpenAI, 2023; Driess et al., 2023; Pope et al., 2023).
Recent efforts have primarily focused on improving the performance of LLM inference on GPUs (Aminabadi et al., 2022; Nvidia, 2022), although GPUs are known for their high power consumption and are less suitable for latency-sensitive workloads (Kim et al., 2023b; Pope et al., 2023). There is also an active body of research dedicated to developing specialized hardware accelerators tailored for Transformer models, with several of these efforts using FPGAs as the target platforms (Hong et al., 2022; Liu et al., 2021; Li et al., 2020a; Peng et al., 2021; Hur et al., 2023; Qi et al., 2021a).
@@
FPGA-based LLM accelerators can be broadly categorized into two architectural paradigms: temporal architecture and spatial architecture. In a temporal architecture, a processing engine (PE) capable of performing various tasks is constructed and reused across different layers and models, as shown in Figure 1(a). For flexibility, these accelerators typically employ an overlay approach (Hong et al., 2022; Li et al., 2020a; Khan et al., 2021), where a virtual hardware architecture that executes instructions is “laid” on top of the physical FPGA fabric. Overlays provide a more restricted configuration space, allowing for quicker compilation with bitstream reuse across multiple models. However, the use of such temporal architecture requires more frequent off-chip memory access, as intermediate results must be written back to memory. This incurs a cost in terms of both latency and energy consumption that is significantly higher than direct on-chip memory access. Additionally, one could argue that an FPGA overlay will inherently be less efficient than its hardened ASIC counterpart.
In contrast, an FPGA-based spatial architecture typically involves the specialization of distinct PEs for specific operators or layers, facilitating direct communication between them using streaming buffers (e.g., FIFOs or multi-buffers) (Xiang et al., 2022; Wang et al., 2021; Umuroglu et al., 2017; Petrica et al., 2020), as depicted in Figure 1(b-c). This dataflow-style execution substantially reduces off-chip memory accesses and enables the concurrent processing of multiple PEs in a pipelined manner. Moreover, the fine-grained programmability of FPGAs allows efficient support of model-specific spatial architectures, which can further leverage efficiency optimizations such as low-bitwidth quantization, custom numerical types, and sparsity (Zhang et al., 2021a; Sun et al., 2022; Yang et al., 2023; Peng et al., 2022). These capabilities can potentially enable highly efficient LLM inference implementations that surpass GPUs, especially in small-batch low-latency scenarios.
@@
Challenge 1: Navigating diverse parallelism in LLMs. The generative inference process of LLMs typically consists of two distinct stages: (1) simultaneously processing user prompts and (2) sequentially generating new tokens in an autoregressive manner. These two stages exhibit significantly different computational and memory characteristics (detailed in §3), making it necessary to tailor hardware accelerators for their specific needs. This challenge cannot be directly addressed by leveraging techniques from the traditional convolutional neural network (CNN) designs (Kim et al., 2023b; Zhang et al., 2015). The large number of parameters and intermediate tensors further complicates the choice between on-chip and off-chip storage. Additionally, harnessing multiple accelerators for distributed LLM inference adds complexity, particularly when dealing with intricate parallelization schemes (Shoeybi et al., 2019; Narayanan et al., 2019; Hong et al., 2022).
Challenge 2: Lack of standard LLM building blocks in hardware accelerators. The rapid evolution of LLM architectures (Bommasani et al., 2021; OpenAI, 2023; Touvron et al., 2023a) contrasts with the comparatively slow pace of hardware development. While a plethora of building blocks for Transformers have been proposed in the software domain (Lefaudeux et al., 2022; Dao et al., 2022; ELS-RD, 2022), the absence of reusable blocks for hardware accelerator design hampers development progress. Many frameworks have been designed to automatically map deep learning models to FPGAs (Zhang et al., 2018; Umuroglu et al., 2017; Fahim et al., 2021; Basalama et al., 2023; Zhang et al., 2020), but they are constrained to small CNN designs and lack support for complicated Transformer models. It is also hard to scale their designs to accommodate large models and multi-die FPGAs.
@@
  • We create a suite of modular and reusable HLS kernels designed for building FPGA-based spatial accelerators for different Transformer models. We plan to open-source this kernel library and expect it to serve as a valuable resource for benchmarking HLS and FPGA acceleration more broadly.
  • Leveraging our kernel library, we design and implement a range of high-performance FPGA-based LLM accelerators that achieve speedups comparable to previous GPU and FPGA-based accelerators. Specifically, for the BERT model, we achieve an 16.1$\times$ speedup over prior FPGA-based accelerators. For GPT generative inference, we achieve speedups of 2.2$\times$ and 1.1$\times$ in prefill and decode stages respectively, when compared to DFX, an FPGA-based overlay architecture. Additionally, our accelerator is 1.9$\times$ faster and 5.7$\times$ more energy-efficient than the A100 GPU in the decode stage.
@@ 2.1. Transformer Models
The Transformer model consists of both encoder and decoder blocks (Vaswani et al., 2017). Recent employment on LLMs mostly uses decoder-only models, which leverage an auto-regressive approach for text generation (Radford et al., 2019; OpenAI, 2023; Touvron et al., 2023a). We will mainly discuss decoder-only models in this paper, but since encoders and decoders share the core building blocks with subtle architectural variances, our approach can also be extended for encoder-only models (Devlin et al., 2018; Lan et al., 2020; Liu et al., 2019).
As illustrated in Figure 2, generative inference of LLMs has two stages: prefill stage and decode stage (Pope et al., 2023). In the prefill stage, the model takes in user prompts, normally a long sequence with $l_{input}$ tokens, goes through the whole Transformer model, and generates the first token. In the decode stage, the model takes in the previously generated token and generates $l_{gen}$ new tokens one at a time in an auto-regressive way. Since each token depends on the previously generated tokens, the decode stage is purely sequential.
We then go through the detailed model architecture. The input tokens are first passed into an embedding layer that maps the discrete tokens into high-dimensional continuous representations while incorporating positional encoding for each token. Subsequently, it generates a tensor (i.e., hidden states) of shape $(l,d)$, where $l$ represents sequence length, and $d$ is the size of hidden dimensions. We omit the batch dimension to simplify the analysis, focusing solely on single-batch inference in this paper, but our approach can be easily extended to different batch sizes for LLM serving (Kwon et al., 2023; Li et al., 2023).
The hidden states then pass through a series of $N$ Transformer blocks. Each Transformer block consists of two sublayers: a multi-head attention (MHA) module and a feed-forward network (FFN). Residual connections and layer normalization (LayerNorm) functions are applied between these sublayers, although the specific order and application may vary across different models (Xiong et al., 2020). The MHA module plays a crucial role in capturing token relationships within the input sequence. The input is initially partitioned into $h$ segments, where $h$ corresponds to the number of attention heads. To compute the attention scores for each head, the input sequence of length $l$ undergoes three linear projections: query, key, and value. These projections, which are trainable, yield matrices $Q$, $K$, and $V$ respectively. Attention scores are then computed using a scaled dot-product (SDP) operation between $Q$, $K$, and $V$, as specified by the formula:
@@
where $d_{k}$ is the size of the hidden dimension. The output from this operation involves $h$ outputs, which are subsequently concatenated and processed through an additional linear projection. In the prefill stage, the generated $K$ and $V$ tensors will be stored as KV cache and later be concatenated before SDP during the decode stage (Pope et al., 2023).
@@
Finally, the output tensor will go through a softmax function to obtain a distribution. The model will sample a token from this distribution and feed it into the decode stage. For encoder-only models, there is only a prefill stage involved, and the distribution will be directly used for different downstream tasks like text classification (Devlin et al., 2018; Liu et al., 2019; Lan et al., 2020).
In this paper, we only focus on analyzing the core Transformer blocks and accelerating them on FPGAs. Embedding layers and output sampling (HuggingFace, 2023; Chen et al., 2023) require extensive random memory accesses, which may not be suitable for FPGA acceleration. Also, they only take a small fraction of overall compute that does not affect the overall latency (Kim et al., 2023b), so we leave them to execute on CPUs or GPUs as usual.
@@ 2.2. Parallelization Schemes
Figure 3. An example of tensor parallelism of a Transformer layer with two devices. TP rank is the unique identifier given to a device within a TP group. SM is the softmax function, and GL is the GeLU function.
As model sizes continue to expand, it becomes increasingly common for a single device to be insufficient for accommodating the entire model. Consequently, the exploration of diverse parallelism schemes within a device and across devices becomes necessary. Parallelism techniques in deep learning can be roughly classified into data, tensor, and pipeline parallelism, together known as 3D parallelism (Shoeybi et al., 2019; Narayanan et al., 2021b; Korthikanti et al., 2022; Chen et al., 2024). During the era of CNNs, data parallelism was the norm, involving the partitioning of input tensors along the batch dimension and their distribution across multiple devices (Li et al., 2020b; Abadi et al., 2016). DeepSpeed ZeRO (Rajbhandari et al., 2020) and FSDP (Meta, 2021) extended data parallelism by proposing a three-stage parallelism strategy that partitions optimizer states, gradients, and parameters to minimize memory usage. However, this approach may incur high communication overheads during inference.
A more recent parallelism scheme for LLM inference is tensor parallelism (TP) (Li et al., 2023; Aminabadi et al., 2022; Shoeybi et al., 2019; Narayanan et al., 2021b; Pope et al., 2023), which distributes model parameters across multiple devices and conducts explicit collective operators to ensure model correctness. Megatron-LM (Shoeybi et al., 2019) is the first to explore tensor parallelism for Transformer-based models, proving to be efficient in both training and inference due to relatively low communication costs. As shown in Figure 3, tensor parallelism requires two all_reduce operators inside a Transformer layer to ensure the results are correct. Our accelerator design also explores tensor parallelism, as detailed in §5.2.
Lastly, pipeline parallelism (Narayanan et al., 2019; Yang et al., 2021; Narayanan et al., 2021a) divides the model across network layers. Multiple layers are grouped into a pipeline stage, and different stages are assigned to different devices. Pipeline parallelism is typically employed across multiple nodes. Since both tensor parallelism and pipeline parallelism handle only portions of the network model, they are collectively referred to as model parallelism. We revisit these parallelism schemes in §5.2.
@@ 3.1. Computational Demands
Our first imperative is to calculate the computational demands of the model. Given that the predominant computation in the Transformer model hinges on general matrix-matrix multiplication (GEMM or Matmul) or general matrix-vector multiplication (GEMV) operations (Pope et al., 2023; Kim et al., 2023b), we employ the number of multiply–accumulates (MACs) as the proxy metric for quantifying compute requirements of the linear layers, as depicted in Table 1.
We denote $X_{(\cdot)}$ as the output tensors from preceding layers, and $W_{(\cdot)}$ represents the weights of corresponding linear layers. For example, $X_{\text{sm}}$ is the output of the softmax operator. Our analysis here is restricted to a single batch; hence the tensors only have two dimensions. We can observe that the computational demand during the prefill stage far surpasses that of the decode stage. In the prefill stage, the required MACs of the two Matmuls within the SDP are quadratic to the sequence length. Consequently, when input sequences exhibit substantial length, attention layers may extend computation time significantly. On the contrary, in the decode stage, each operator processes a single token at a time, making the MACs independent of the sequence length except for SDP.
@@
  Linear Layer Abbreviations Computation Prefill Decode

### 07 Apr 2024 v2

@@ Abstract.
To enable more productive implementations of an LLM model on FPGAs, we further provide a library of high-level synthesis (HLS) kernels that are composable and reusable. This library will be made available as open-source. To validate the effectiveness of both our analytical model and HLS library, we have implemented BERT and GPT2 on an AMD Xilinx Alveo U280 FPGA device. Experimental results demonstrate our approach can achieve up to 13.4$\times$ speedup when compared to previous FPGA-based accelerators for the BERT model. For GPT generative inference, we attain a 2.2$\times$ speedup compared to DFX, an FPGA overlay, in the prefill stage, while achieving a 1.9$\times$ speedup and a 5.7$\times$ improvement in energy efficiency compared to the NVIDIA A100 GPU in the decode stage.
FPGA, high-level synthesis, large language models, hardware acceleration
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: TRETS’24; May 5–8, 2024; Orlando, FLisbn: 978-1-4503-XXXX-X/18/06ccs: Hardware Hardware-software codesignccs: Computing methodologies Neural networks
@@ 1. Introduction
The rapid advancement of Transformer-based large language models (LLMs) (vaswani2017transformer, ; rishi2021foundation, ) has sparked a revolution across a wide range of natural language processing tasks, such as conversational AI (openai2023gpt4, ; zheng2023chatbotarena, ; cohen2022lamda, ) and code generation (chen2021codex, ; li2022alphacode, ; nijkamp2023codegen, ). Recent research has brought to light the phenomenon of “emergence” in LLMs, where advanced capabilities become evident as the models scale up to billions of parameters (wei2022emergence, ; wei2022cot, ). However, supporting this unprecedented scale poses significant challenges, particularly in terms of computational and memory resources. At the same time, the increasing use of LLMs in interactive applications like voice assistants and autonomous systems requires hardware accelerators capable of providing both low latency and high energy efficiency (openai2023gpt4, ; driess2023palme, ; pope2022googleinf, ).
Recent efforts have primarily focused on improving the performance of LLM inference on GPUs (aminabadi2022dsinf, ; fastertransformer2022, ), although GPUs are known for their high power consumption and are less suitable for latency-sensitive workloads (kim2023survey, ; pope2022googleinf, ). There is also an active body of research dedicated to developing specialized hardware accelerators tailored for Transformer models, with several of these efforts using FPGAs as the target platforms (hong2022dfx, ; liu2021fqbert, ; li2020ftrans, ; peng2021cbbp, ; hur2023flexrun, ; qi2021iccad, ).
@@
FPGA-based LLM accelerators can be broadly categorized into two architectural paradigms: temporal architecture and spatial architecture. In a temporal architecture, a processing engine (PE) capable of performing various tasks is constructed and reused across different layers and models, as shown in Figure 1(a). For flexibility, these accelerators typically employ an overlay approach (hong2022dfx, ; li2020ftrans, ; khan2021npe, ), where a virtual hardware architecture that executes instructions is “laid” on top of the physical FPGA fabric. Overlays provide a more restricted configuration space, allowing for quicker compilation with bitstream reuse across multiple models. However, the use of such temporal architecture requires more frequent off-chip memory access, as intermediate results must be written back to memory. This incurs a cost in terms of both latency and energy consumption that is significantly higher than direct on-chip memory access. Additionally, one could argue that an FPGA overlay will inherently be less efficient than its hardened ASIC counterpart.
In contrast, an FPGA-based spatial architecture typically involves the specialization of distinct PEs for specific operators or layers, facilitating direct communication between them using streaming buffers (e.g., FIFOs or multi-buffers) (xiang2022heteroflow, ; wang2021autosa, ; yaman2017finn, ; petrica2020dataflowcnn, ), as depicted in Figure 1(b-c). This dataflow-style execution substantially reduces off-chip memory accesses and enables the concurrent processing of multiple PEs in a pipelined manner. Moreover, the fine-grained programmability of FPGAs allows efficient support of model-specific spatial architectures, which can further leverage efficiency optimizations such as low-bitwidth quantization, custom numerical types, and sparsity (zhang2021fracbnn, ; sun2022autovit, ; yang2023aim, ; peng2022sparsefpga, ). These capabilities can potentially enable highly efficient LLM inference implementations that surpass GPUs, especially in small-batch low-latency scenarios.
@@
Challenge 1: Navigating diverse parallelism in LLMs. The generative inference process of LLMs typically consists of two distinct stages: (1) simultaneously processing user prompts and (2) sequentially generating new tokens in an autoregressive manner. These two stages exhibit significantly different computational and memory characteristics (detailed in §3), making it necessary to tailor hardware accelerators for their specific needs. This challenge cannot be directly addressed by leveraging techniques from the traditional convolutional neural network (CNN) designs (kim2023survey, ; zhang2015cnnfpga, ). The large number of parameters and intermediate tensors further complicates the choice between on-chip and off-chip storage. Additionally, harnessing multiple accelerators for distributed LLM inference adds complexity, particularly when dealing with intricate parallelization schemes (shoeybi2019megatron, ; narayanan2019pipedream, ; hong2022dfx, ).
Challenge 2: Lack of standard LLM building blocks in hardware accelerators. The rapid evolution of LLM architectures (rishi2021foundation, ; openai2023gpt4, ; touvron2023llama, ) contrasts with the comparatively slow pace of hardware development. While a plethora of building blocks for Transformers have been proposed in the software domain (xFormers2022, ; dao2022flashattention, ; kernl2022, ), the absence of reusable blocks for hardware accelerator design hampers development progress. Many frameworks have been designed to automatically map deep learning models to FPGAs (zhang2018dnnbuilder, ; yaman2017finn, ; fahim2021hls4ml, ; suhail2023flexcnn, ; zhang2020dnnexplorer, ), but they are constrained to small CNN designs and lack support for complicated Transformer models. It is also hard to scale their designs to accommodate large models and multi-die FPGAs.
@@
  • We create a suite of modular and reusable HLS kernels designed for building FPGA-based spatial accelerators for different Transformer models. We plan to open-source this kernel library [^1] https://github.com/cornell-zhang/allo/tree/main/examples and expect it to serve as a valuable resource for benchmarking HLS and FPGA acceleration more broadly.
  • Leveraging our kernel library, we design and implement a range of high-performance FPGA-based LLM accelerators that achieve speedups comparable to previous GPU and FPGA-based accelerators. Specifically, for the BERT model, we achieve a 13.4$\times$ speedup over prior FPGA-based accelerators. For GPT generative inference, we achieve speedups of 2.2$\times$ and 1.1$\times$ in prefill and decode stages respectively, when compared to DFX, an FPGA-based overlay architecture. Additionally, our accelerator is 1.9$\times$ faster and 5.7$\times$ more energy-efficient than the A100 GPU in the decode stage.
@@ 2.1. Transformer Models
The Transformer model consists of both encoder and decoder blocks (vaswani2017transformer, ). Recent employment on LLMs mostly uses decoder-only models, which leverage an auto-regressive approach for text generation (radford2019gpt2, ; openai2023gpt4, ; touvron2023llama, ). We will mainly discuss decoder-only models in this paper, but since encoders and decoders share the core building blocks with subtle architectural variances, our approach can also be extended for encoder-only models (devlin2018bert, ; lan2020albert, ; liu2019roberta, ).
As illustrated in Figure 2, generative inference of LLMs has two stages: prefill stage and decode stage (pope2022googleinf, ). In the prefill stage, the model takes in user prompts, normally a long sequence with $l_{\text{input}}$ tokens, goes through the whole Transformer model, and generates the first token. In the decode stage, the model takes in the previously generated token and generates $l_{\text{gen}}$ new tokens one at a time in an auto-regressive way. Since each token depends on the previously generated tokens, the decode stage is purely sequential.
We then go through the detailed model architecture. The input tokens are first passed into an embedding layer that maps the discrete tokens into high-dimensional continuous representations while incorporating positional encoding for each token. Subsequently, it generates a tensor (i.e., hidden states) of shape $(l,d)$, where $l$ represents sequence length, and $d$ is the size of hidden dimensions. We omit the batch dimension to simplify the analysis, focusing solely on single-batch inference in this paper, but our approach can be easily extended to different batch sizes for LLM serving by adding an additional batch dimension (kwon2023vllm, ; li2023alpaserve, ).
The hidden states then pass through a series of $N$ Transformer blocks. Each Transformer block consists of two sublayers: a multi-head attention (MHA) module and a feed-forward network (FFN). Residual connections and layer normalization (LayerNorm) functions are applied between these sublayers, although the specific order and application may vary across different models (xiong2020preln, ). The MHA module plays a crucial role in capturing token relationships within the input sequence. The input is initially partitioned into $h$ segments, where $h$ corresponds to the number of attention heads. To compute the attention scores for each head, the input sequence of length $l$ undergoes three linear projections: query, key, and value. These projections, which are trainable, yield matrices $Q$, $K$, and $V$ respectively. Attention scores are then computed using a scaled dot-product (SDP) operator between $Q$, $K$, and $V$, as specified by the formula:
@@
where $d_{k}$ is the size of the hidden dimension. The output from this operator involves $h$ outputs, which are subsequently concatenated and processed through an additional linear projection. In the prefill stage, the generated $K$ and $V$ tensors will be stored as KV cache and later be concatenated before SDP during the decode stage (pope2022googleinf, ).
@@
Finally, the output tensor will go through a softmax function to obtain a distribution. The model will sample a token from this distribution and feed it into the decode stage. For encoder-only models, there is only a prefill stage involved, and the distribution will be directly used for different downstream tasks like text classification (devlin2018bert, ; liu2019roberta, ; lan2020albert, ).
In this paper, we only focus on analyzing the core Transformer blocks and accelerating them on FPGAs. Embedding layers and output sampling (generation_strategies, ; chen2023speculative, ) require extensive random memory accesses, which may not be suitable for FPGA acceleration. Also, they only take a small fraction of overall compute that does not affect the overall latency (kim2023survey, ), so we leave them to execute on CPUs or GPUs as usual.
@@ 2.2. Parallelization Schemes
Figure 3. An example of tensor parallelism of a Transformer layer with two devices. TP rank is the unique identifier given to a device within a TP group. SM is the softmax function, LN is LayerNorm, and GL is the GeLU function.
As model sizes continue to expand, it becomes increasingly common for a single device to be insufficient for accommodating the entire model. Consequently, the exploration of diverse parallelization schemes within a device and across devices becomes necessary. Parallelism techniques in deep learning can be roughly classified into data, tensor, and pipeline parallelism, together known as 3D parallelism (shoeybi2019megatron, ; narayanan2021megatronv2, ; korthikanti2022reducing, ; chen2024slapo, ). During the era of CNNs, data parallelism was the norm, involving the partitioning of input tensors along the batch dimension and their distribution across multiple devices (li2020ptddp, ; abadi2016tensorflow, ). DeepSpeed ZeRO (rajbhandari2020zero, ) and FSDP (fsdp, ) extended data parallelism by proposing a three-stage parallelism strategy that partitions optimizer states, gradients, and parameters to minimize memory usage. However, this approach may incur high communication overheads during inference.
A more recent parallelism scheme for LLM inference is tensor parallelism (TP) (li2023alpaserve, ; aminabadi2022dsinf, ; shoeybi2019megatron, ; narayanan2021megatronv2, ; pope2022googleinf, ), which distributes model parameters across multiple devices and conducts explicit collective operations to ensure model correctness. Megatron-LM (shoeybi2019megatron, ) is the first to explore tensor parallelism for Transformer-based models, proving to be efficient in both training and inference due to relatively low communication costs. As shown in Figure 3, tensor parallelism requires two all_reduce operations inside a Transformer layer to ensure the results are correct. Our accelerator design also explores tensor parallelism, as detailed in §3.4.2.
Lastly, pipeline parallelism (narayanan2019pipedream, ; yang2021pipemare, ; narayanan2021pipe2bw, ) divides the model across network layers. Multiple layers are grouped into a pipeline stage, and different stages are assigned to different devices. Pipeline parallelism is typically employed across multiple nodes. Since both tensor parallelism and pipeline parallelism handle only portions of the network model, they are collectively referred to as model parallelism. We revisit these parallelization schemes in §3.4.2.
@@ 3.1. Computational Demands
Our first imperative is to calculate the computational demands of the model. Given that the predominant computation in the Transformer model is general matrix-matrix multiplication (GEMM or Matmul) or general matrix-vector multiplication (GEMV) (pope2022googleinf, ; kim2023survey, ), we employ the number of multiply–accumulates (MACs) as the proxy metric for quantifying compute requirements of the linear layers, as depicted in Table 1. For non-linear layers such as softmax and GeLU functions, they are elementwise operators that can be easily fused with the GEMM kernels in a pipeline design without affecting the final performance. More experimental results are provided in §6.3.
We denote $X_{(\cdot)}$ as the output tensors from preceding layers, and $W_{(\cdot)}$ represents the weights of corresponding linear layers. For example, $X_{\text{sm}}$ is the output of the softmax operator. Our analysis here is restricted to a single batch; hence the tensors only have two dimensions. We can observe that the computational demand during the prefill stage far surpasses that of the decode stage. In the prefill stage, the required MACs of the two Matmuls within the SDP are quadratic to the sequence length (i.e., $l^{2}d$). Consequently, when input sequences exhibit substantial length, attention layers may extend computation time significantly. On the contrary, in the decode stage, each operator processes a single token at a time, making the MACs independent of the sequence length except for SDP.
@@
  Linear Layer Abbreviations Input Matrices Prefill Decode